OpenStudy (anonymous):

What is the value of z in the equation 5z − 9 = 36? 4 5 6 9

OpenStudy (anonymous):

@texaschic101

OpenStudy (welshfella):

OK first add 9 to each side of the equation 5z - 9 = 36 +9 +9 what do we get?

OpenStudy (anonymous):

i have no idea

OpenStudy (welshfella):

- 9 + 9 = 0 right?

OpenStudy (anonymous):

yea

OpenStudy (welshfella):

and 36 + 9 = what?

OpenStudy (anonymous):

45

OpenStudy (welshfella):

right so we have 5z on left side and 45 on the right side:- 5z = 45 we added 9 so we wanted 5z on its own on left side

OpenStudy (welshfella):

* as we wanted

OpenStudy (welshfella):

ok?

OpenStudy (anonymous):

im still confused lol please dontget mad

OpenStudy (anonymous):

oh wait i think i get it so would it be 40

OpenStudy (welshfella):

5z - 9 = 36 first we aim to get 5z on its own to do this we add 9 to both sides ( -9+9 = 0 so we are left with 5z on the left. now so that the equation still balances we have also to add 9 to the right side.

OpenStudy (anonymous):

yeah

OpenStudy (welshfella):

its a bit like a chemical balance - if you put an extra weight on one pan , to make it balance again , you have to put the same weight on the other pan.

OpenStudy (anonymous):

oh i figured it out its 9 thanks si much il give u a medal

OpenStudy (welshfella):

yes you are right z = 45/5 = 9

OpenStudy (welshfella):

good job

OpenStudy (welshfella):